OpenStudy (aravindg):

ax/cos theta)+(by/sin theta)=a^2-b^2 and (axsin theta/cos^2 theta)-(bycos theta/sin^2 theta)=0 show that (ax)^2/3 +(by)^2/3=(a^2-b^2)^2/3

OpenStudy (anonymous):

From first equation \begin{eqnarray*} \frac{ax}{\cos\theta}+\frac{by}{\sin\theta} & = & a^2-b^2 \\ \frac{ax\sin\theta+by\cos\theta}{\cos\theta\sin\theta} & = & a^2-b^2 \\ ax\sin\theta+by\cos\theta & = & (a^2-b^2)\cos\theta\sin\theta \end{eqnarray*} From second equation \begin{eqnarray*} \frac{ax\sin\theta}{\cos^2\theta}-\frac{by\cos\theta}{\sin^2\theta} & = & 0 \\ \frac{ax\sin^3\theta-by\cos^3\theta}{\cos^2\theta\sin^2\theta} & = & 0 \\ ax\sin^3\theta-by\cos^3\theta & = & 0 \\ ax\sin^3\theta & = & by\cos^3\theta \\ ax\tan^3\theta & = & by \end{eqnarray*} Substituting in first equation: \begin{eqnarray*} ax\sin\theta+by\cos\theta & = & (a^2-b^2)\cos\theta\sin\theta \\ ax\sin\theta+(ax\tan^3\theta)\cos\theta & = & (a^2-b^2)\cos\theta\sin\theta \\ ax\sin\theta+ax\tan^2\theta\frac{\sin\theta}{\cos\theta}\cos\theta & = & (a^2-b^2)\cos\theta\sin\theta \\ ax\sin\theta+ax\tan^2\theta\sin\theta & = & (a^2-b^2)\cos\theta\sin\theta \\ ax+ax\tan^2\theta & = & (a^2-b^2)\cos\theta \\ ax(1+\tan^2\theta) & = & (a^2-b^2)\cos\theta \\ ax\sec^2\theta & = & (a^2-b^2)\cos\theta \\ ax & = & (a^2-b^2)\cos^3\theta \end{eqnarray*} ... to be continued.

OpenStudy (aravindg):

continue plzzz

OpenStudy (nilankshi):

hmmmmm

OpenStudy (anonymous):

Hence \[(ax)^\frac{2}{3} = (a^2-b^2)^\frac{2}{3}\cos^2\theta\] and \begin{eqnarray*} (by)^\frac{2}{3} & = & (ax)^\frac{2}{3}\tan^2\theta \\ & = & ((a^2-b^2)^\frac{2}{3}\cos^2\theta)\tan^2\theta \\ & = & (a^2-b^2)^\frac{2}{3}\cos^2\theta\frac{\sin^2\theta}{\cos^2\theta} \\ (by)^\frac{2}{3} & = & (a^2-b^2)^\frac{2}{3}\sin^2\theta \end{eqnarray*} So finally, \begin{eqnarray*} (ax)^\frac{2}{3}+(by)^\frac{2}{3} & = & (a^2-b^2)^\frac{2}{3}\cos^2\theta + (a^2-b^2)^\frac{2}{3}\sin^2\theta\\ & = & (a^2-b^2)^\frac{2}{3}(\cos^2\theta+\sin^2\theta) \\ (ax)^\frac{2}{3}+(by)^\frac{2}{3} & = & (a^2-b^2)^\frac{2}{3} \end{eqnarray*} as required.

OpenStudy (aravindg):

man is this much steps necessary is there an easier way???

OpenStudy (anonymous):

Well I tried to make it simpler but I couldn't... :(

OpenStudy (aravindg):

wel thx i hav some moe questions plz dont go

OpenStudy (aravindg):

k?

OpenStudy (anonymous):

Have about 45 more mins here.

OpenStudy (aravindg):

wow thx

OpenStudy (nilankshi):

u ar in which grade

OpenStudy (aravindg):

if sin (theta+A)/sin(theta+B)=sqrt(sin 2A/sin2B),prove that tan^2theta=tan( A)*tan (B)

OpenStudy (aravindg):

ME??

OpenStudy (nilankshi):

yes

OpenStudy (aravindg):

11th

OpenStudy (aravindg):

y u ask??

OpenStudy (nilankshi):

i am not able to understand it

OpenStudy (nilankshi):

i am in 9th thats why

OpenStudy (aravindg):

I asked why u asked my grade??

OpenStudy (aravindg):

u a girl??

OpenStudy (nilankshi):

yes

OpenStudy (aravindg):

got too many fans lol

OpenStudy (aravindg):

xactxx do help me

OpenStudy (nilankshi):

hmmmmm

OpenStudy (aravindg):

:)

OpenStudy (nilankshi):

u have also many fans

OpenStudy (nilankshi):

i am your fan also

OpenStudy (aravindg):

thnk u

OpenStudy (nilankshi):

welcome

OpenStudy (aravindg):

XACTXXX HELPPPPPP

OpenStudy (anonymous):

hold on I first write down the answer on a piece of paper

OpenStudy (aravindg):

2.If alpha and beta be 2 distinct roots satisfying equation a cos theta+bsin theta=c,Show that cos (alpha+beta)=(a^2-b^2)/(a^2+b^2)

OpenStudy (aravindg):

............

OpenStudy (aravindg):

plz answer both and questions over

OpenStudy (nilankshi):

hmmmmm

OpenStudy (anonymous):

ok I got the second one I think \begin{eqnarray*} \frac{\sin(\theta+A)}{\sin(\theta+B)} & = &\sqrt{\frac{\sin 2A}{\sin 2B}} \\ \Big(\frac{\sin\theta\cos A+\cos\theta\sin A}{\sin\theta\cos B+\cos\theta\sin B}\Big)^2 & = & \frac{2\sin A\cos A}{2\sin B\cos B} \\ (\sin\theta\cos A+\cos\theta\sin A)^2\sin B\cos B & = & (\sin\theta\cos B+\cos\theta\sin B)^2\sin A\cos A \\ (\sin^2\theta\cos^2 A+2\cos\theta\sin\theta\cos A\sin A+\cos^2\theta\sin^2 A)\sin B\cos B & = & (\sin^2\theta\cos^2 B+2\cos\theta\sin\theta\cos B\sin B+\cos^2\theta\sin^2 B)\sin A\cos A \\ \sin^2\theta\cos^2 A\sin B\cos B+2\cos\theta\sin\theta\cos A\sin A\sin B\cos B+\cos^2\theta\sin^2 A\sin B\cos B & = & \sin^2\theta\cos^2 B\sin A\cos A+2\cos\theta\sin\theta\cos B\sin B\sin A\cos A+\cos^2\theta\sin^2 B\sin A\cos A \end{eqnarray*} \begin{eqnarray*} \sin^2\theta\cos^2 A\sin B\cos B+\cos^2\theta\sin^2 A\sin B\cos B & = & \sin^2\theta\cos^2 B\sin A\cos A+\cos^2\theta\sin^2 B\sin A\cos A \\ \sin^2\theta\cos^2 A\sin B\cos B-\sin^2\theta\cos^2 B\sin A\cos A & = & \cos^2\theta\sin^2 B\sin A\cos A-\cos^2\theta\sin^2 A\sin B\cos B \\ \sin^2\theta\cos A\cos B(\cos A\sin B-\cos B\sin A) & = & \cos^2\theta\sin A\sin B(\cos A\sin B-cos B\sin A) \\ \sin^2\theta\cos A\cos B & = & \cos^2\theta\sin A\sin B \\ \frac{\sin^2\theta}{\cos^2\theta} & = & \frac{\sin A\sin B}{\cos A\cos B} \\ \tan^2\theta & = & \tan A\tan B \end{eqnarray*} as required.

OpenStudy (aravindg):

can u read???

OpenStudy (anonymous):

right click that and click 'show source'. It will help

OpenStudy (aravindg):

its not understandable

OpenStudy (aravindg):

this show source

OpenStudy (anonymous):

\[\tiny\tiny\tiny\tiny\tiny\tiny\small\small\small\small\small{\begin{eqnarray*} \frac{\sin(\theta+A)}{\sin(\theta+B)} & = &\sqrt{\frac{\sin 2A}{\sin 2B}} \\ \Big(\frac{\sin\theta\cos A+\cos\theta\sin A}{\sin\theta\cos B+\cos\theta\sin B}\Big)^2 & = & \frac{2\sin A\cos A}{2\sin B\cos B} \\ (\sin\theta\cos A+\cos\theta\sin A)^2\sin B\cos B & = & (\sin\theta\cos B+\cos\theta\sin B)^2\sin A\cos A \\ (\sin^2\theta\cos^2 A+2\cos\theta\sin\theta\cos A\sin A+\cos^2\theta\sin^2 A)\sin B\cos B & = & (\sin^2\theta\cos^2 B+2\cos\theta\sin\theta\cos B\sin B+\cos^2\theta\sin^2 B)\sin A\cos A \\ \sin^2\theta\cos^2 A\sin B\cos B+2\cos\theta\sin\theta\cos A\sin A\sin B\cos B+\cos^2\theta\sin^2 A\sin B\cos B & = & \sin^2\theta\cos^2 B\sin A\cos A+2\cos\theta\sin\theta\cos B\sin B\sin A\cos A+\cos^2\theta\sin^2 B\sin A\cos A \end{eqnarray*}}\] I've tried my best

OpenStudy (anonymous):

AravindG, right click then choose Zoom Trigger, Click, then click on the mathematical text. You could scroll through the solution afterwards.

OpenStudy (aravindg):

still right part is missing